# On the Optimal Rate of Convergence for Translation-Invariant 1D Quantum Walks

### v1

@@ Abstract.
We study the convergence rate of translation-invariant discrete-time quantum dynamics on a one-dimensional lattice. We prove that the cumulative distributions function of the ballistically scaled position ${\mathds{X}}({n})/{n}$ after $n$ steps converges at a rate of ${n}^{-1/3}$ in the Lévy metric as ${n}\to\infty$. In the special case of step-coin quantum walks with two-dimensional coin space, we recover the same convergence rate for the supremum distance and prove optimality.
@@ 1. Introduction
Quantum walks have since their introduction [ABN+01] attracted a lot of attention as an experimentally realizable platform with many applications, e.g., for search algorithms in quantum computing [Amb07, LCE+10, Por13]. Particular emphasis has been put on their analogy to the classical random walk, where the main computational advantage of the quantum walk is the ballistic propagation of information. Especially, analogous results to the classical central limit theorem have by now been established in various settings with versatile mathematical approaches, see for example [Kon02, GJS04, Kon05, AVWW11, ST12, Suz16, Wad20] and references therein.
@@
In comparison to the classical random walk, which has a convergence order of $n^{-1/2}$ as the number of steps $n$ goes to $\infty$, we in fact discover that quantum walks converge slower at the order of $n^{-1/3}$. This is due to the fact that most information is located directly in the ballistically propagating wavefront region, where the convergence is thus slowed down. In view of the many applications of quantum walks in quantum computing and their experimental realization, our results will provide an important ingredient in estimating computational errors in the future. Natural extensions of our result would also be to study non translation-invariant walks or higher-dimensional systems.
@@
We further prove a corresponding upper bound for general 1D lattice discrete time quantum dynamics for the Lévy metric, which is in fact an important intermediate step in the proof of the sup-norm bound as well.
Our proof essentially consists of two parts: First, we prove an adapted Esseen inequality, estimating differences of cumulative distribution functions from the corresponding characteristic functions. Applying it to the setting of 1D quantum walks yields the convergence rate $n^{-1/3}$ for the Lévy distance. We then carefully analyze the concrete position distribution for the step-coin quantum walk, by building on results from [ST12]. In this second step, we especially need to focus on a treatment of the wavefront region mentioned above, which eventually yields the optimal convergence rate in the supremum distance.
@@
In Section˜2.1, we present the exact setting of this article and state our main results Theorems˜2.1 and 2.2. In Section˜3, we then prove a generalized Esseen–Zolotarev type inequality, which applied to our setting yields the proof of Theorem˜2.1. In the final Section˜4, we then sharpen the estimates for coin-step quantum walks to prove Theorem˜2.2.
@@ 2.1.1. General Setup
We are here most interested in choosing $A={\mathds{X}}$ the (selfadjoint) position operator given by $({\mathds{X}}\psi)_{k}=k\psi_{k}$, $\psi\in{\mathcal{H}}$. Then for given initial state $\rho$, we analyze the ballistically rescaled position CDF. After $n$ steps, it is given by $F_{{\mathds{X}}/n}^{W^{n}\rho W^{-n}}=F^{\rho}_{{\mathds{X}}_{n}}$ with ${\mathds{X}}_{n}\coloneqq W^{-n}{\mathds{X}}W^{n}/n$.
@@ 2.1.2. Translation-Invariance and Weak Limit Theorems
We will restrict our attention to translation-invariant time step operators $W$, i.e., assume that $W$ commutes with the right shift operator $(T\psi)_{k}\coloneqq\psi_{k-1}$. Denoting by ${\mathcal{F}}:\ell^{2}({\mathbb{Z}};{\mathcal{K}})\to L^{2}({\mathds{T}};{\mathcal{K}})$ with ${\mathds{T}}=[0,2\pi)$ the usual unitary Fourier transform defined by
$({\mathcal{F}}\psi)(p)=\frac{1}{\sqrt{2\pi}}\sum_{k\in{\mathbb{Z}}}\psi_{k}e^{ikp},\quad p\in{\mathds{T}},$
translation-invariance is equivalent to the existence of a map $\widehat{W}:{\mathds{T}}\to{\mathcal{U}}({\mathcal{K}})$ such that
@@
We will here further assume that $\widehat{W}(p)$ has purely discrete spectrum for all $p\in{\mathds{T}}$, i.e., there exists a (finite or countable) family of orthogonal projections $(\Pi_{k}(p))_{k\in{\mathcal{I}}}$ on $\mathcal{K}$} and a corresponding family of real numbers $(\omega_{k}(p))_{k\in{\mathcal{I}}}\subset{\mathbb{R}}$ such that
@@
If $\omega_{k}$ is differentiable in almost every $p\in{\mathds{T}}$, we can then define the so-called velocity operator as the selfadjoint operator given by
@@
In this case the strong convergence $e^{it{\mathds{X}}_{n}}\xrightarrow{n\to\infty}e^{itV}$ for all $t\in{\mathbb{R}}$ is well-established as a type of weak central limit theorem for quantum walks [Kon02, GJS04, AVWW11, Suz16]. By the Portmanteau theorem, this is equivalent to the convergence
$F^{\rho}_{{\mathds{X}}_{n}}(x)\xrightarrow{n\to\infty}F^{\rho}_{V}(x)$
@@ 2.2. Lévy Metric and General Upper Bound
In ˜2.2 additionally assume that $\omega_{k}\in C^{2}({\mathds{T}};{\mathbb{R}})$ and $\Pi_{k}\in C^{1}({\mathds{T}};{\mathcal{B}}({\mathcal{H}}))$ such that
$\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathds{T}}}\|\Pi_{k}^{\prime}(p)\|_{{\mathcal{B}}({\mathcal{K}})}<\infty\quad\mbox{and}\quad\sup_{k\in{\mathcal{I}},\,p\in{\mathds{T}}}|\omega_{k}^{\prime\prime}(p)|<\infty.$
Then, for any density matrix $\rho$ satisfying $\operatorname{tr}(\lvert{\mathds{X}}\rvert\rho)<\infty$, there exists a constant $C>0$ such that
$L(F_{{\mathds{X}}_{n}}^{\rho},F^{\rho}_{V})\leq Cn^{-\frac{1}{3}}.$
@@ 2.3. Kolmogorov Metric and Optimality for Step-Coin Walks
Let us now introduce the concept of the usual step-coin quantum walk with the two-dimensional coin space ${\mathcal{K}}={\mathbb{C}}^{2}$. In this setting, recalling the right shift operator $T$ from above and using the natural identification ${\mathcal{H}}=\ell^{2}({\mathbb{Z}};{\mathbb{C}}^{2})\cong\ell^{2}({\mathbb{Z}})\oplus\ell^{2}({\mathbb{Z}})$, the step operator $S$ is defined as
@@
Note that ${\mathcal{F}}S{\mathcal{F}}^{*}(p)=e^{ip}\oplus e^{-ip}$, i.e., ${\mathcal{F}}S{\mathcal{F}}^{*}$ is the direct sum of multiplication operators. We will then call $W$ the time-step operator of a step-coin walk on $\mathcal{H}$} if $W=SC$ for some $C\in{\mathcal{U}}({\mathbb{C}}^{2})$ acting as $(C\psi)_{k}=C\psi_{k}$, $k\in{\mathbb{Z}}$ called coin operator.
@@
$C_{1}n^{-\frac{1}{3}}\leq\bigl\|F_{{\mathds{X}}_{n}}^{\rho}-F_{V}^{\rho}\bigr\|_{\infty}\leq C_{2}n^{-\frac{1}{3}}.$
@@
In contrast to the classical Berry–Esseen rate, which is $n^{-1/2}$, the one-dimensional step-coin quantum walk converges strictly slower. The reduced rate arises from the formation of ballistic wavefronts: near the edges of the support of its density, $F_{V}^{\rho}$ develops a nondifferentiable cusp and the approximation error is dominated by this boundary behaviour. Analyzing this boundary region is the main ingredient to the proof of the above theorem.
@@
It would be desirable to also obtain a corresponding lower bound for the Lévy metric in the coin-step quantum walk case. Our method in this case yields that for all $\varepsilon>0$
$n^{-\frac{2}{3}-\varepsilon}\lesssim L(F_{{\mathds{X}}_{n}^{\rho}},F_{V}^{\rho})\lesssim n^{-\frac{1}{3}},$
@@ 3.2. Estimates on The Characteristic Functions
$\widehat{F}_{{\mathds{X}}_{n}}^{\rho}(\lambda)=\operatorname{tr}\big(\rho W^{-n}e^{i\lambda{\mathds{X}}}W^{n}\big),\qquad\widehat{F}_{V}^{\rho}(\lambda)=\operatorname{tr}\big(\rho e^{i\lambda V}\big).$
@@
Let $W\in{\mathcal{U}}({\mathcal{H}})$ satisfy the assumptions of Theorem˜2.1 and let $\rho\in{\mathcal{T}}({\mathcal{H}})$ be a density matrix satisfying $\operatorname{tr}(\lvert{\mathds{X}}\rvert\rho)<\infty$. Then, for all $n\in{\mathbb{N}}$ and $\lambda\in{\mathbb{R}}$,
$\left\lvert\widehat{F}_{{\mathds{X}}_{n}}^{\rho}\left(\frac{\lambda}{n}\right)-\widehat{F}_{V}^{\rho}(\lambda)\right\rvert\leq\frac{\lvert\lambda\rvert^{2}}{n}\sup_{k\in{\mathcal{I}},p\in{\mathds{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\frac{\lvert\lambda\rvert}{n}\left(\operatorname{tr}(\lvert{\mathds{X}}\rvert\rho)+\left(\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathds{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert\right)\right).$
@@
$\left\lvert\widehat{F}_{{\mathds{X}}_{n}}^{\rho}\left(\frac{\lambda}{n}\right)-\widehat{F}_{V}^{\rho}(\lambda)\right\rvert\leq\left\lvert\operatorname{tr}\left(\rho W^{-n}e^{i\frac{\lambda}{n}\mathds{X}}W^{n}\left(1-e^{-i\frac{\lambda}{n}\mathds{X}}\right)\right)\right\rvert+\left\lvert\operatorname{tr}\left(\rho\left(e^{i\lambda V}-W^{-n}e^{i\frac{\lambda}{n}\mathds{X}}W^{n}e^{-i\frac{\lambda}{n}\mathds{X}}\right)\right)\right\rvert$.
@@
$\left\lvert\operatorname{tr}\left({W^{-n}}e^{i\frac{\lambda}{n}\mathds{X}}W^{n}\rho-{W^{-n}}e^{i\frac{\lambda}{n}\mathds{X}}W^{n}e^{-i\frac{\lambda}{n}{\mathds{X}}}\rho\right)\right\rvert\leq\frac{\lvert\lambda\rvert}{n}\operatorname{tr}\left(\lvert\mathds{X}\rvert\rho\right),$
@@
Since $W^{-n}e^{i\frac{\lambda}{n}\mathds{X}}W^{n}$ is unitary, we have
$\left\lvert\operatorname{tr}\left({W^{-n}}e^{i\frac{\lambda}{n}\mathds{X}}W^{n}\rho-{W^{-n}}e^{i\frac{\lambda}{n}\mathds{X}}W^{n}e^{-i\frac{\lambda}{n}\mathds{X}}\rho\right)\right\rvert\leq\left\lvert\operatorname{tr}\left((1-e^{i\frac{\lambda}{n}\mathds{X}})\rho\right)\right\rvert$.
@@
$\left\lvert\operatorname{tr}\left(\big(1-e^{i\frac{\lambda}{n}\mathds{X}}\big)\rho\right)\right\rvert$ $\leq\sum_{i}\sum_{x\in{\mathbb{Z}}}\lambda_{i}\left\lvert 1-e^{i\frac{\lambda}{n}x}\right\rvert\|\psi_{i}(x)\|_{\mathcal{K}}^{2}\leq\sum_{i}\sum_{x\in{\mathbb{Z}}}\lambda_{i}\left\lvert\frac{\lambda}{n}x\right\rvert\|\psi_{i}(x)\|_{\mathcal{K}}^{2}=\frac{\lvert\lambda\rvert}{n}\operatorname{tr}\left(\lvert\mathds{X}\rvert\rho\right).\qed$
@@
Let $\rho\in{\mathcal{T}}({\mathcal{H}})$ be a density matrix, and $W\in{\mathcal{U}}({\mathcal{H}})$ be translation-invariant with locally purely discrete spectrum, i.e., ˜2.2 holds, and assume that $\omega_{k}\in C^{2}({\mathds{T}};{\mathbb{R}})$ and $\Pi_{k}\in C^{1}({\mathds{T}};{\mathcal{B}}({\mathcal{H}}))$. Further, define the asymptotic velocity operator by ˜2.3 as before. Then, for any $n\in{\mathbb{N}}$ and $\lambda\in{\mathbb{R}}$,
$\left\lvert\operatorname{tr}\left(\rho\left(e^{i\lambda V}-W^{-n}e^{i\frac{\lambda}{n}\mathds{X}}W^{n}e^{-i\frac{\lambda}{n}\mathds{X}}\right)\right)\right\rvert\leq\frac{\lvert\lambda\rvert^{2}}{n}\sup_{p\in{\mathds{T}},k\in{\mathcal{I}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\frac{\lvert\lambda\rvert}{n}\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathds{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{{\mathcal{B}}({\mathcal{K}})},$
@@
$W^{-n}e^{i\frac{\lambda}{n}\mathds{X}}W^{n}e^{-i\frac{\lambda}{n}\mathds{X}}\xrightarrow{n\to\infty}e^{i\lambda V}$
with respect to the operator norm. We note that by assumption, both operators in the difference act locally after Fourier transforming, i.e., for $p\in{\mathds{T}}$, we have as operator identities on $\mathcal{K}$}
${\mathcal{F}}W^{-n}e^{i\frac{\lambda}{n}\mathds{X}}W^{n}e^{-i\frac{\lambda}{n}\mathds{X}}{\mathcal{F}}^{*}(p)=\sum_{k,\ell\in{\mathcal{I}}}e^{in(\omega_{\ell}(p+\tfrac{\lambda}{n})-\omega_{k}(p))}\Pi_{k}(p)\Pi_{\ell}\left(p+\frac{\lambda}{n}\right),$
@@
(3.5) $\big\|W^{-n}e^{i\frac{\lambda}{n}{\mathds{X}}}W^{n}e^{-i\frac{\lambda}{n}{\mathds{X}}}-U_{n}\big\|\leq\frac{\lvert\lambda\rvert}{n}\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathds{T}}}\left\lVert\Pi_{k}^{\prime}(p)\right\rVert_{{\mathcal{B}}(\mathcal{K})}$
(3.6) $\left\lVert e^{i\lambda V}-U_{n}\right\rVert\leq\frac{\lvert\lambda\rvert^{2}}{n}\sup_{k\in{\mathcal{I}},p\in{\mathds{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert$.
Note that for a local operator $T$ on $L^{2}({\mathds{T}};{\mathcal{K}})$, i.e., an operator acting as $Tf(p)=T(p)f(p)$ with some $T(p)\in{\mathcal{B}}({\mathcal{K}})$, is equivalent to estimating the operator norm pointwise, namely,
(3.7) $\|T\|_{{\mathcal{B}}(L^{2}({\mathds{T}};{\mathcal{K}}))}=\operatorname*{ess\,sup}_{p\in{\mathds{T}}}\|T(p)\|_{{\mathcal{B}}({\mathcal{K}})}.$
By unitarity of the Fourier transform, we can thus estimate the defining expression pointwise in $\mathds{T}$}. More precisely, fixing, $p\in{\mathds{T}}$, we find
$\big\lVert\mathcal{F}\big({W}^{-n}e^{i\frac{\lambda}{n}\mathds{X}}W^{n}e^{-i\frac{\lambda}{n}\mathds{X}}-U_{n}\big)\mathcal{F}^{*}(p)\big\rVert_{{\mathcal{B}}({\mathcal{K}})}=\left\lVert\sum_{k,\ell\in{\mathcal{I}}}e^{in\left(\omega_{\ell}(p+\frac{\lambda}{n})-\omega_{k}(p)\right)}\Pi_{k}(p)\big(\Pi_{\ell}(p+\tfrac{\lambda}{n})-\Pi_{\ell}(p)\big)\right\rVert_{{\mathcal{B}}({\mathcal{K}})}$
@@
$\qquad\qquad\leq\frac{\lvert\lambda\rvert}{n}\sum_{\ell\in{\mathcal{I}}}\sup_{p\in{\mathds{T}}}\left\lVert\Pi_{\ell}^{\prime}(p)\right\rVert_{{\mathcal{B}}({\mathcal{K}})},$

### v2

@@ Abstract.
We study the convergence rate of translation-invariant discrete-time quantum dynamics on a one-dimensional lattice. We prove that the cumulative distributions function of the ballistically scaled position X(n)/nX({n})/{n} after $n$ steps converges at a rate of ${n}^{-1/3}$ in the Lévy metric as ${n}\to\infty$. In the special case of shift-coin quantum walks with two-dimensional coin space, we recover the same convergence rate for the supremum distance and prove optimality.
@@ 1. Introduction
Quantum walks have since their introduction [ABN+01] attracted a lot of attention as an experimentally realizable platform [PLP+08, KFC+09, SCP+10, GAS+13] with many applications, e.g., for search algorithms in quantum computing [Amb07, LCE+10, Por13, Mon16] or quantum simulation [AGW12, Chi09]. Particular emphasis has been put on the analogy of quantum walks to the classical random walk, where the main computational advantage of the quantum walk is the ballistic propagation of information. Especially, analogous results to the classical central limit theorem have by now been established in various settings with versatile mathematical approaches, see for example [Kon02, GJS04, Kon05, AVWW11, ST12, Suz16, Wad20] and references therein.
@@
In comparison to the classical random walk, which has a convergence order of $n^{-1/2}$ as the number of steps $n$ goes to $\infty$, we in fact discover that quantum walks converge slower at the order of $n^{-1/3}$. This is due to the fact that most information is located directly in the ballistically propagating wavefront region, where the convergence is thus slowed down. In view of the many applications of quantum walks in quantum computing and their experimental realization, our results will provide an important ingredient in estimating computational errors in the future. Natural extensions of our result would also be to study non translation-invariant walks or higher-dimensional systems. As far as one-dimensional systems go, our upper bound on the convergence rate in fact holds for a large class of unitary lattice dynamics, especially including the so-called split-step quantum walks, which can be described in the framework of CMV matrices [BHJ03, CMV03, CGMV12, CF24].
@@
We further prove a corresponding upper bound for general discrete-time quantum dynamics on the one-dimensional lattice in the Lévy metric, which is in fact an important intermediate step in the proof of the sup-norm bound as well.
Our proof essentially consists of two parts: First, we prove an adapted Esseen inequality, estimating differences of cumulative distribution functions from the corresponding characteristic functions. Applying it to the setting of 1D quantum walks yields the convergence rate $n^{-1/3}$ for the Lévy distance. We then carefully analyze the concrete position distribution for the shift-coin quantum walk, by building on results from [ST12]. In this second step, we especially need to focus on a treatment of the wavefront region mentioned above, which eventually yields the optimal convergence rate in the supremum distance.
@@
In Section˜2.1, we present the exact setting of this article and state our main results Theorems˜2.1 and 2.2. In Section˜3, we then prove a generalized Esseen–Zolotarev type inequality, which applied to our setting yields the proof of Theorem˜2.1. In the final Section˜4, we then sharpen the estimates for shift-coin quantum walks to prove Theorem˜2.2.
@@ 2.1.1. General Setup
We are here most interested in choosing $A=X$ the (selfadjoint) position operator given by $(X\psi)_{k}=k\psi_{k}$, $\psi\in{\mathcal{H}}$. Then for given initial state $\rho$, we analyze the ballistically rescaled position CDF. After $n$ steps, it is given by $F_{X/n}^{W^{n}\rho W^{-n}}=F^{\rho}_{X_{n}}$ with $X_{n}\coloneqq W^{-n}XW^{n}/n$.
@@ 2.1.2. Translation-Invariance and Weak Limit Theorems
We will restrict our attention to translation-invariant time step operators $W$, i.e., assume that $W$ commutes with the right shift operator $(T\psi)_{k}\coloneqq\psi_{k-1}$. Denoting by ${\mathcal{F}}:\ell^{2}({\mathbb{Z}};{\mathcal{K}})\to L^{2}({\mathbb{T}};{\mathcal{K}})$ with ${\mathbb{T}}=[0,2\pi)$ the usual unitary Fourier transform defined by
$({\mathcal{F}}\psi)(p)=\frac{1}{\sqrt{2\pi}}\sum_{k\in{\mathbb{Z}}}\psi_{k}e^{ikp},\quad p\in{\mathbb{T}},$
translation-invariance is equivalent to the existence of a map $\widehat{W}:{\mathbb{T}}\to{\mathcal{U}}({\mathcal{K}})$ such that
@@
We will here further assume that $\widehat{W}(p)$ has purely discrete spectrum for all $p\in{\mathbb{T}}$, i.e., there exists a (finite or countable) family of orthogonal projections $(\Pi_{k}(p))_{k\in{\mathcal{I}}}$ on $\mathcal{K}$} and a corresponding family of real numbers $(\omega_{k}(p))_{k\in{\mathcal{I}}}\subset{\mathbb{R}}$ such that
@@
If $\omega_{k}$ is differentiable in almost every $p\in{\mathbb{T}}$, we can then define the so-called velocity operator as the selfadjoint operator given by
@@
In this case the strong convergence $e^{itX_{n}}\xrightarrow{n\to\infty}e^{itV}$ for all $t\in{\mathbb{R}}$ is well-established as a type of weak central limit theorem for quantum walks [Kon02, GJS04, AVWW11, Suz16]. By the Portmanteau theorem, this is equivalent to the convergence
$F^{\rho}_{X_{n}}(x)\xrightarrow{n\to\infty}F^{\rho}_{V}(x)$
@@ 2.2. Lévy Metric and General Upper Bound
In ˜2.2 additionally assume that $\omega_{k}\in C^{2}({\mathbb{T}};{\mathbb{R}})$ and $\Pi_{k}\in C^{1}({\mathbb{T}};{\mathcal{B}}({\mathcal{H}}))$ such that
$\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathbb{T}}}\|\Pi_{k}^{\prime}(p)\|_{{\mathcal{B}}({\mathcal{K}})}<\infty\quad\mbox{and}\quad\sup_{k\in{\mathcal{I}},\,p\in{\mathbb{T}}}|\omega_{k}^{\prime\prime}(p)|<\infty.$
Then, for any density matrix $\rho$ satisfying $\operatorname{tr}(\lvert X\rvert\rho)<\infty$, there exists a constant $C>0$ such that
$L(F_{X_{n}}^{\rho},F^{\rho}_{V})\leq Cn^{-\frac{1}{3}}.$
@@ 2.3. Kolmogorov Metric and Optimality for Step-Coin Walks
Let us now introduce the concept of the usual shift-coin quantum walk with the two-dimensional coin space ${\mathcal{K}}={\mathbb{C}}^{2}$. In this setting, recalling the right shift operator $T$ from above and using the natural identification ${\mathcal{H}}=\ell^{2}({\mathbb{Z}};{\mathbb{C}}^{2})\cong\ell^{2}({\mathbb{Z}})\oplus\ell^{2}({\mathbb{Z}})$, the step operator $S$ is defined as
@@
Note that ${\mathcal{F}}S{\mathcal{F}}^{*}(p)=e^{ip}\oplus e^{-ip}$, i.e., ${\mathcal{F}}S{\mathcal{F}}^{*}$ is the direct sum of multiplication operators. We will then call $W$ the time-step operator of a shift-coin walk on $\mathcal{H}$} if $W=SC$ for some $C\in{\mathcal{U}}({\mathbb{C}}^{2})$ acting as $(C\psi)_{k}=C\psi_{k}$, $k\in{\mathbb{Z}}$ called coin operator.
@@
$C_{1}n^{-\frac{1}{3}}\leq\bigl\|F_{X_{n}}^{\rho}-F_{V}^{\rho}\bigr\|_{\infty}\leq C_{2}n^{-\frac{1}{3}}.$
@@
In contrast to the classical Berry–Esseen rate, which is $n^{-1/2}$, the one-dimensional shift-coin quantum walk converges strictly slower. The reduced rate arises from the formation of ballistic wavefronts: near the edges of the support of its density, $F_{V}^{\rho}$ develops a nondifferentiable cusp and the approximation error is dominated by this boundary behaviour. Analyzing this boundary region is the main ingredient to the proof of the above theorem.
@@
It would be desirable to also obtain a corresponding lower bound for the Lévy metric in the shift-coin quantum walk case. Our method in this case yields that for all $\varepsilon>0$
$n^{-\frac{2}{3}-\varepsilon}\lesssim L(F_{X_{n}^{\rho}},F_{V}^{\rho})\lesssim n^{-\frac{1}{3}},$
@@ 3.2. Estimates on The Characteristic Functions
$\widehat{F}_{X_{n}}^{\rho}(\lambda)=\operatorname{tr}\big(\rho W^{-n}e^{i\lambda X}W^{n}\big),\qquad\widehat{F}_{V}^{\rho}(\lambda)=\operatorname{tr}\big(\rho e^{i\lambda V}\big).$
@@
Let $W\in{\mathcal{U}}({\mathcal{H}})$ satisfy the assumptions of Theorem˜2.1 and let $\rho\in{\mathcal{T}}({\mathcal{H}})$ be a density matrix satisfying $\operatorname{tr}(\lvert X\rvert\rho)<\infty$. Then, for all $n\in{\mathbb{N}}$ and $\lambda\in{\mathbb{R}}$,
$\left\lvert\widehat{F}_{X_{n}}^{\rho}\left(\frac{\lambda}{n}\right)-\widehat{F}_{V}^{\rho}(\lambda)\right\rvert\leq\frac{\lvert\lambda\rvert^{2}}{n}\sup_{k\in{\mathcal{I}},p\in{\mathbb{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\frac{\lvert\lambda\rvert}{n}\left(\operatorname{tr}(\lvert X\rvert\rho)+\left(\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathbb{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert\right)\right).$
@@
$\left\lvert\widehat{F}_{X_{n}}^{\rho}\left(\frac{\lambda}{n}\right)-\widehat{F}_{V}^{\rho}(\lambda)\right\rvert\leq\left\lvert\operatorname{tr}\left(\rho W^{-n}e^{i\frac{\lambda}{n}{X}}W^{n}\left(1-e^{-i\frac{\lambda}{n}{X}}\right)\right)\right\rvert+\left\lvert\operatorname{tr}\left(\rho\left(e^{i\lambda V}-W^{-n}e^{i\frac{\lambda}{n}{X}}W^{n}e^{-i\frac{\lambda}{n}{X}}\right)\right)\right\rvert$.
@@
$\left\lvert\operatorname{tr}\left({W^{-n}}e^{i\frac{\lambda}{n}{X}}W^{n}\rho-{W^{-n}}e^{i\frac{\lambda}{n}{X}}W^{n}e^{-i\frac{\lambda}{n}X}\rho\right)\right\rvert\leq\frac{\lvert\lambda\rvert}{n}\operatorname{tr}\left(\lvert{X}\rvert\rho\right),$
@@
Since $W^{-n}e^{i\frac{\lambda}{n}{X}}W^{n}$ is unitary, we have
$\left\lvert\operatorname{tr}\left({W^{-n}}e^{i\frac{\lambda}{n}{X}}W^{n}\rho-{W^{-n}}e^{i\frac{\lambda}{n}{X}}W^{n}e^{-i\frac{\lambda}{n}{X}}\rho\right)\right\rvert\leq\left\lvert\operatorname{tr}\left((1-e^{i\frac{\lambda}{n}{X}})\rho\right)\right\rvert$.
@@
$\left\lvert\operatorname{tr}\left(\big(1-e^{i\frac{\lambda}{n}{X}}\big)\rho\right)\right\rvert$ $\leq\sum_{i}\sum_{x\in{\mathbb{Z}}}\lambda_{i}\left\lvert 1-e^{i\frac{\lambda}{n}x}\right\rvert\|\psi_{i}(x)\|_{\mathcal{K}}^{2}\leq\sum_{i}\sum_{x\in{\mathbb{Z}}}\lambda_{i}\left\lvert\frac{\lambda}{n}x\right\rvert\|\psi_{i}(x)\|_{\mathcal{K}}^{2}=\frac{\lvert\lambda\rvert}{n}\operatorname{tr}\left(\lvert{X}\rvert\rho\right).\qed$
@@
Let $\rho\in{\mathcal{T}}({\mathcal{H}})$ be a density matrix, and $W\in{\mathcal{U}}({\mathcal{H}})$ be translation-invariant with locally purely discrete spectrum, i.e., ˜2.2 holds, and assume that $\omega_{k}\in C^{2}({\mathbb{T}};{\mathbb{R}})$ and $\Pi_{k}\in C^{1}({\mathbb{T}};{\mathcal{B}}({\mathcal{H}}))$. Further, define the asymptotic velocity operator by ˜2.3 as before. Then, for any $n\in{\mathbb{N}}$ and $\lambda\in{\mathbb{R}}$,
$\left\lvert\operatorname{tr}\left(\rho\left(e^{i\lambda V}-W^{-n}e^{i\frac{\lambda}{n}{X}}W^{n}e^{-i\frac{\lambda}{n}{X}}\right)\right)\right\rvert\leq\frac{\lvert\lambda\rvert^{2}}{n}\sup_{p\in{\mathbb{T}},k\in{\mathcal{I}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert+\frac{\lvert\lambda\rvert}{n}\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathbb{T}}}\lVert\Pi_{k}^{\prime}(p)\rVert_{{\mathcal{B}}({\mathcal{K}})},$
@@
$W^{-n}e^{i\frac{\lambda}{n}{X}}W^{n}e^{-i\frac{\lambda}{n}{X}}\xrightarrow{n\to\infty}e^{i\lambda V}$
with respect to the operator norm. We note that by assumption, both operators in the difference act locally after Fourier transforming, i.e., for $p\in{\mathbb{T}}$, we have as operator identities on $\mathcal{K}$}
${\mathcal{F}}W^{-n}e^{i\frac{\lambda}{n}{X}}W^{n}e^{-i\frac{\lambda}{n}{X}}{\mathcal{F}}^{*}(p)=\sum_{k,\ell\in{\mathcal{I}}}e^{in(\omega_{\ell}(p+\tfrac{\lambda}{n})-\omega_{k}(p))}\Pi_{k}(p)\Pi_{\ell}\left(p+\frac{\lambda}{n}\right),$
@@
(3.5) $\big\|W^{-n}e^{i\frac{\lambda}{n}X}W^{n}e^{-i\frac{\lambda}{n}X}-U_{n}\big\|\leq\frac{\lvert\lambda\rvert}{n}\sum_{k\in{\mathcal{I}}}\sup_{p\in{\mathbb{T}}}\left\lVert\Pi_{k}^{\prime}(p)\right\rVert_{{\mathcal{B}}(\mathcal{K})}$
(3.6) $\left\lVert e^{i\lambda V}-U_{n}\right\rVert\leq\frac{\lvert\lambda\rvert^{2}}{n}\sup_{k\in{\mathcal{I}},p\in{\mathbb{T}}}\lvert\omega_{k}^{\prime\prime}(p)\rvert$.
Note that for a local operator $T$ on $L^{2}({\mathbb{T}};{\mathcal{K}})$, i.e., an operator acting as $Tf(p)=T(p)f(p)$ with some $T(p)\in{\mathcal{B}}({\mathcal{K}})$, is equivalent to estimating the operator norm pointwise, namely,
(3.7) $\|T\|_{{\mathcal{B}}(L^{2}({\mathbb{T}};{\mathcal{K}}))}=\operatorname*{ess\,sup}_{p\in{\mathbb{T}}}\|T(p)\|_{{\mathcal{B}}({\mathcal{K}})}.$
By unitarity of the Fourier transform, we can thus estimate the defining expression pointwise in $\mathbb{T}$}. More precisely, fixing, $p\in{\mathbb{T}}$, we find
$\big\lVert\mathcal{F}\big({W}^{-n}e^{i\frac{\lambda}{n}{X}}W^{n}e^{-i\frac{\lambda}{n}{X}}-U_{n}\big)\mathcal{F}^{*}(p)\big\rVert_{{\mathcal{B}}({\mathcal{K}})}=\left\lVert\sum_{k,\ell\in{\mathcal{I}}}e^{in\left(\omega_{\ell}(p+\frac{\lambda}{n})-\omega_{k}(p)\right)}\Pi_{k}(p)\big(\Pi_{\ell}(p+\tfrac{\lambda}{n})-\Pi_{\ell}(p)\big)\right\rVert_{{\mathcal{B}}({\mathcal{K}})}$
@@
$\qquad\qquad\leq\frac{\lvert\lambda\rvert}{n}\sum_{\ell\in{\mathcal{I}}}\sup_{p\in{\mathbb{T}}}\left\lVert\Pi_{\ell}^{\prime}(p)\right\rVert_{{\mathcal{B}}({\mathcal{K}})},$